A) If you want to be 95​% confident of estimating the population mean to within a sampling error of plus or minus 2 and the standard deviation is assumed to be 16​, what sample size is​ required?

The sample size required is _____

​(Round up to the nearest​ integer.)  

B) If you want to be 95​% confident of estimating the population proportion to within a sampling error of plus or minus 0.06, what sample size is​ needed?

A sample size of ____ is needed.

​(Round up to the nearest​ integer.)  
Answer #1

Solution :

Given that,

A) standard deviation = \sigma = 16

margin of error = E = 2

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Sample size = n = ((Z\alpha/2 * \sigma ) / E)2

= ((1.96 * 16 ) / 2)2

= 245.8

Sample size = 24

B)  \hat p = 1 - \hat p = 0.5

margin of error = E = 0.06

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.96 / 0.06)2 * 0.5 * 0.5

= 266.7

sample size = 267